OpenStudy (anonymous):

I need help like quick please Rhett decides to build a square room for his movie and music collection. If the area of the room is 4x2 + 28x + 49 square feet, what is the length of one side of the room? (7x + 2) feet (2x + 7) feet (2x − 7) feet (7x − 2) feet and Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____ Numerical Answers Expected!

OpenStudy (anonymous):

4x^2+28x+49 = (2x+7)^2 Since the room is square, its area formula is A = s^2, where s = length of one side so the length of one side = the square root of the equation, which = (2x+7) feet

OpenStudy (anonymous):

@Shawna.marie

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

No problem! :)

OpenStudy (reemii):

the formula is \((a+b)^2=a^2+2ab + b^2\). you have \((x+*)^2 = x^2 + 2x* + *^2\). You are told that \(2x* = 16x\). You can find \(*\) and then the missing term.

OpenStudy (reemii):

edit: \(2x* = -16x\) (minus sign)